OpenStudy (anonymous):

PLEASE HELP The function f(t) = 5 cos(pi over 4t) + 11 represents the tide in Dark Sea. It has a maximum of 16 feet when time (t) is 0 and a minimum of 6 feet. The sea repeats this cycle every 8 hours. After six hours, how high is the tide?

OpenStudy (anonymous):

Ive typed it into my calculator several different ways but am unable to get any of these answers A. 11 feet B. 6 feet C. 8.5 feet D. 10.5 feet

OpenStudy (anonymous):

I keep getting 15.9

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@robtobey

jimthompson5910 (jim_thompson5910):

Plug in t = 6 and evaluate. f(t) = 5*cos((pi/4)*t) + 11 f(6) = 5*cos((pi/4)*6) + 11 f(6) = 5*cos(3pi/2) + 11 f(6) = 5*0 + 11 ... use the unit circle here f(6) = 0 + 11 f(6) = 11 So at t = 6 hours, the height is 11 feet

OpenStudy (anonymous):

Thanks so much omgg...